OpenStudy (anonymous):

What is the sum of an 8–term geometric series if the first term is –11, the last term is 859,375, and the common ratio is –5? –143,231 716,144 144,177 –36,047

OpenStudy (anonymous):

716,144?

OpenStudy (anonymous):

I got -716144

OpenStudy (anonymous):

what formula you used? \[S=a\frac{1-r^n}{1-r}\] this?

OpenStudy (anonymous):

No it looks like: a1 - a1 (r)^n Sn = _______________ 1 - r

OpenStudy (anonymous):

it's same

OpenStudy (anonymous):

a1-a1r^n=a1(1-r^n)

OpenStudy (anonymous):

i calculated wrongly, you are correct, sorry

OpenStudy (anonymous):

Haha it's ok, thank you

OpenStudy (anonymous):

i calculated as \[S=a\frac{r^n-1}{1-r}\] mixed r^n and 1 ><

OpenStudy (anonymous):

Ty!!! This was correct!